OpenStudy (anonymous):

simplify (x^(-3)y^(-1))^-1

jimthompson5910 (jim_thompson5910):

Multiply the outer exponent by all the inner exponents to get x^((-3)*(-1))y^((-1)*(-1)) = x^3y^1 = x^3y So the answer is x^3y

OpenStudy (anonymous):

no nut thnx anyway

OpenStudy (anonymous):

but*

jimthompson5910 (jim_thompson5910):

what's wrong?

OpenStudy (anonymous):

x isnt bein set to one whole exponent its two expressions each with their own radical multiplied by each other and the entire expression is being put to the -1 power

jimthompson5910 (jim_thompson5910):

can you draw it out

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

yeah that's what I thought it was originally \[\Large (x^{-3}y^{-1})^{-1}=x^3y\]

OpenStudy (anonymous):

ok thnx thought u meant x^(3y)

jimthompson5910 (jim_thompson5910):

oh, no sry

OpenStudy (anonymous):

lol k thnx

jimthompson5910 (jim_thompson5910):

probably should have just drawn it out lol